OpenStudy (anonymous):

Can you help me with this problem? Find the equation of a line with the given points. (5, 4) (5, 6)

OpenStudy (anonymous):

My answer is y = -x + 10

OpenStudy (anonymous):

first find the slope and then use point slope formula to find an equation

OpenStudy (anonymous):

y-y1=m(x-x1)

OpenStudy (anonymous):

where y1 and x1 are your x and y coordinates and m is your slope

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

The equation for the slope is x = 5 and y is any real number

OpenStudy (anonymous):

That confuses me.

OpenStudy (anonymous):

no don't look for the slope, because both first coordinates are the same that makes the slope undefined, so forget about finding it. since the first coordinate in both points is 5, that makes the equation of the line \[x=5\]

OpenStudy (anonymous):

if both of the second coordinates are the same, for example if you see \[(2,7),(5,7)\] the equation of the line would be \(y=7\)

OpenStudy (anonymous):

ok, that makes sense.